What is the overall change in internal energy for System A after both of the following processes have occurred?

Process 1: System A absorbs 10 kJ of heat from the surroundings and does 8 kJ of work on its surroundings.

Process 2: System A releases 5 kJ of heat into the surroundings and 10000 J of work is done on the system.

  1. 2 kJ, ΔE = negative
  2. 7 kJ, ΔE = positive
  3. 33 kJ, ΔE = negative
  4. 9982 J, ΔE = positive
  5. 3000 J, ΔE = negative
Answer #1

From first law of thermodynamics.

DE = q+w

Process 1:

w = work is done by the system = - 8 KJ

Heat absrobed by the system = 10 KJ

The change in internal energy of the system, DE = 10-8 = 2 KJ

Process 2:

w = work is done on the system = 10000 J = 10 KJ

Heat released by the system = -5 KJ

The change in internal energy of the system, DE = 10-5 = 5 KJ

But, Internal energy is an extensive property and is additive in nature.

Therefore, the overall change in internal energy for System A after both of the following processes have occurred = 5 + 2 = 7 KJ and is positive.